# Also at ] The European Space Agency Centre for Earth Observation, Frascati, Italy

### v4

@@ Abstract
The spontaneous emergence of coherent structures from chaotic backgrounds is a hallmark of active biological swarms. We investigate this self-organization by simulating an ensemble of polar chiral active agents that couple locally via a Kuramoto interaction. We demonstrate that the system’s transition from chaos to active turbulence is characterized by quantized loop phase currents and coherent clustering, and that this transition is strictly governed by a kinetic Turing instability. By deriving the continuum kinetic theory for the model, we identify that the competition between local phase-locking and active agent motility selects a critical structural wavenumber. The instability drives the system into a state of developed turbulence that exhibits stable, robust power-laws in spectral density, suggestive of universality and consistent with observations from a broad range of turbulent phenomena. Our results bridge the gap between discrete chimera states and continuous fluid turbulence, suggesting that the statistical laws of active matter can arise from fundamental kinetic instability criteria.
@@ II The model
Conversely, our model is a spatial ensemble of active chiral agents, swarming in a periodic bounded 2D box, representing active matter turbulence. The Kuramoto coupling term itself first-order; attraction and inhibition take place through the curious interplay between spatial structure and the agents that constitute this structure. We must therefore derive a kinetic extension of the Turing instability into active matter. By so doing, we will bridge the Kuramoto model with plasma physics (Landau damping) and swarming (Vicsek model).
@@ III Kinetic Turing instability
We shall insert the standard Kuramoto phase-locking dynamic into the kinetic Vlasov-Fokker-Planck equation as the reaction-term, preserving the core Kuramoto physics within a continuum field description. We linearize this equation, which describes long-range interactions in many-body systems, i.e. kinetic plasma physics, and the ansatz that we shall use to linearize this equation is identical to the one used to derive Landau damping in plasmas [25]. Of some importance, chiral agent frequency, or frustration $\omega_{i}$, mimics thermal anisotropies in kinetic plasmas, while $G$ mimics electric fields. At the same time, the physical transport of phase information involved in the Kuramoto-coupling means that we are explicitly deriving a kinetic instability threshold for Vicsek-swarming and chiral active matter [8, 9, 10], and explicit connections to plasma physics remain a future endeavor.
@@
follows from Eq. (2). Eq. (5) describes the phase velocity, with the complex field $Z$ mediating the Gaussian kernel interaction (Eq. 3), now expressed as,
@@
$D\frac{\partial^{2}}{\partial\theta^{2}}(e^{in\theta})=-D(in)^{2}e^{in\theta}=\\ =-D(-n^{2})e^{in\theta}=+Dn^{2}e^{in\theta},$ (12)
@@
The analogue to Landau damping in plasma physics then lies in the breakdown of damping for a particular wavenumber, by the competition between $\hat{G}(k)$ and $\chi(k)$ in Eq. (23).
@@
We have implemented a numerical Monte Carlo integration [27] to solve Eqs. (23) and (24) using the finite set of oscillator frequencies from the distribution of intrinsic agent frustration $\omega_{i}$ in Figure 1, which we show in Figure 2a). We observe a minimum for the value $A_{0}\approx 0.2$, for a critical wavenumber of $k_{crit}\approx 0.86$. We note a relatively shallow minimum around $A_{crit}$, meaning that a broad range of wavenumbers will be amplified. In Figure 2b), we show the result of the linear stability analysis applied to all frustration distributions ranging from “slow” to “fast” in Figure 1, demonstrating that global synchronization ($k_{crit}\to 0$ is achieved for distribution multipliers lower than around $10$, at which point global synchronization causes a uniform drift in the agent ensemble. Critical couplings strengths then increase logarithmically with frequency distribution $\omega_{i}$.
@@ IV Results
The wavelength associated with the Turing-instability analysis (Figure 2a)) is indicated with dashed, black line in the first column of Figure 3, and we observe that it coincides with the dominant wavenumber in the structure, providing clear empirical evidence for the analytical derivations in Section III.
The fast-oscillating modes of Figure 3 start in a chaotic, symmetric state (top row). Symmetry breaks with the triggering of the kinetic Turing instability, and the system is organized in islands with coherent phase evolution, forming local, stable vortices. These vortices must be understood in terms of the quantized loop currents predicted by Ref. [33], characterized by the sum of phase differences around closed loop in the network being equal to an integer multiple of $2\pi$. Figure 4 demonstrates this, showing the instantaneous velocity of some 2000 agents undergoing clustering. We observe that tracing a circle around the vortices will cycle through all the phases from 0 to $2\pi$, with a topological winding number of -1 (a counter-clockwise vortex) [33]. We also note that (1) the general motion of guiding centers is perpendicular to the motion of the oscillators, and (2) the attraction (guiding centers) is considerably slower than the inhibition (agent motility), providing additional justification for the derived Turing instability.
The characteristic spatial organization that we see in Figures 3 and 4 are robust in the simulations of fast-mode oscillators, suggesting that the formation of quantized loop phase currents in ensembles of coupled oscillators may be a universal quality in coupled chiral active matter subject to synchronization rules.
@@
Figure 5 demonstrates that the emergence of structure in the agent ensemble reliably occur for base coupling strengths that exceed the critical value $A_{crit}\approx 0.06$ for the kinetic Turing instability, determined after solving Eqs. (23, 24) for an $\omega_{i}$ distribution multiplier of $100$. This is heralded by the power spectral density of the spatial distribution of active agents (red) and guiding centers (blue) to settle on a power law behavior with values roughly between -1.5 and -3.
@@ V Discussion
We have simulated ensembles of polar chiral active matter agents that are subject to phase synchronization through a Kuramoto-Sakaguchi-like phase interaction, demonstrating that complex structures emerge from a kinetic Turing instability triggered in the relatively simple interactions between locally coupled oscillators. Through the rigorous derivation of a continuum field theory [28] and empirical evidence (Figure 5), we have demonstrated the crucial role played by the kinetic Turing instability within the chiral active matter ensemble. This configuration builds on the noise-multiplicative self-organization framework of Ref. [3] and validates the theoretical foundation by Ref. [33] in that stable, quantized loop currents of phase information emerge in the ensemble.
Unlike Ref. [17], where the prerequisites for a Turing instability in a fixed oscillator network were coded into a higher-order Kuramoto interaction, we have demonstrated that a kinetic Turing Turing instability can trigger in the spatial distribution of chiral active matter. Our model relies on a standard localized Kuramoto interaction, where the attractive force mediated by the kernel $G$ creates a positive feedback loop that aligns phase-synchronized agents. Opposing this, the agents’ swim speed $v_{0}$ coupled with their intrinsic frequency dispersion $\omega_{i}$ acts as an effective diffusion mechanism, transporting phase information away from ordered regions and breaking down order. This creates a competition between local activation and kinetic inhibition, driving the bifurcation observed in the linear stability analysis and triggering the instability.
@@
A crucial ingredient in our model is the heterogenity in $\omega_{i}$ [3], the powerlaw distribution in Figure 1. We have confirmed, through rigorous testing, that varying the shape (powerlaw) of this distribution does not substantially change the simulation’s empirical outcome, and we have verified that the stability analysis is not sensitive to which particular powerlaw we select. Nevertheless, it remains for future efforts to map out the specific mechanism governing the non-linear transfer of energy in the system. In this regard, a related paper, Ref. [34], derives the hydrodynamic limit of our model, and rigorously demonstrates that widening the distribution in $\omega_{i}$ injects sufficient enstrophy to drive an inertial cascade [3] that eventualy culminates in the emergence of inviscid Euler turbulence in the chiral active matter model.
@@
Our model is, however, constricted to a class of active polar chiral matter, and though the derivation in Section III explicitly bridge kinetic plasma physics and active matter turbulence modeling, future efforts should explicitly address whether self-organized turbulence can enter into the present reductionist explanations for geophysical and astrophysical turbulent structuring [3].
@@ VI Conclusion
We have modeled chiral active matter turbulence by coupling self-propelled agents locally with a Kuramoto-like interaction, endowing the agents with an intrinsic and heterogenic frustration, following Ref. [3]. We have simulated the ensemble, demonstrating that stable spectral scaling laws emerge in the resulting active matter clustering. To explain the results, we derived the spatial pattern formation rules in the active matter, identifying the coupling threshold as a kinetic Turing instability [28]. Following the kinetic theory framework established for active fluids [51, 52, 42], we analytically determined the critical wavenumber selected by the competition between Kuramoto phase-locking and active chiral transport. This provides a rigorous microscopic derivation for the emergent ’chimera-like’ patterns often studied in discrete topologies [17, 24], unifying them under a single continuum instability criterion.

### v5

@@ Abstract
The spontaneous emergence of coherent structures from chaotic backgrounds is a hallmark of active biological swarms. We investigate this self-organization by simulating an ensemble of polar chiral active agents that couple locally via a Kuramoto interaction. We demonstrate that the system’s transition from chaos to active turbulence is characterized by quantized loop phase currents and coherent clustering, and that this transition is strictly governed by a kinetic Turing instability. By deriving the continuum kinetic theory for the model, we identify that the competition between local phase-locking and active agent motility selects a critical structural wavenumber. The instability then drives the system into a state of developed, active turbulence that exhibits stable, robust power-laws in spectral density, suggestive of universality and consistent with observations from a broad range of turbulent phenomena. Our results bridge the gap between discrete chimera states and continuous fluid turbulence, suggesting that the statistical scaling laws of active turbulence can arise from fundamental kinetic instability criteria.
@@ II The model
Conversely, our model is a spatial ensemble of active chiral agents, swarming in a periodic bounded 2D box, representing active matter turbulence. The Kuramoto coupling term itself is conventional; attraction and inhibition take place through the curious interplay between spatial structure and the agents that constitute this structure. We must therefore derive a kinetic extension of the Turing instability into active matter. By so doing, we will bridge the Kuramoto model with plasma physics (Landau damping).
@@ III Kinetic Turing instability
We shall insert the standard Kuramoto phase-locking dynamic into the kinetic Vlasov-Fokker-Planck equation as the reaction-term, preserving the core Kuramoto physics within a continuum field description. We linearize this equation, which describes long-range interactions in many-body systems, i.e. kinetic plasmas, and the ansatz that we shall use to linearize this equation is identical to the one used to derive Landau damping in plasmas [25, 26]. Of some importance, chiral agent frequency, or frustration $\omega_{i}$, mimics thermal anisotropy in kinetic plasmas, while $G$ mimics electric fields. At the same time, the physical transport of phase information involved in the Kuramoto-coupling means that we are explicitly deriving a kinetic instability threshold for Vicsek-swarming and chiral active matter [8, 9, 10], and direct a application to plasma physics remain a future endeavor.
@@
follows from the deterministic part of Eq. (2), after expanding the $\sin$ term. Eq. (5) describes the phase velocity, with the complex field $Z$ mediating the Gaussian kernel interaction (Eq. 3), now expressed as,
@@
$D\frac{\partial^{2}}{\partial\theta^{2}}(e^{in\theta})=Dn^{2}e^{in\theta},$ (12)
@@
The apt analogue to Landau damping in plasma physics then lies in the breakdown of damping for a particular wavenumber, by the competition between $\hat{G}(k)$ and $\chi(k)$ in Eq. (23).
@@
We have implemented a numerical Monte Carlo integration [27] to solve Eqs. (23) and (24) using the oscillator frequencies from the distribution of intrinsic agent frustration $\omega_{i}$ in Figure 1, which we show in Figure 2a). We observe a minimum for the value $A_{0}\approx 0.2$, for a critical wavenumber of $k_{crit}\approx 0.86$. We note a relatively shallow minimum around $A_{crit}$, meaning that a broad range of wavenumbers will be amplified. In Figure 2b), we show the result of the linear stability analysis applied to all frustration distributions ranging from “slow” to “fast” in Figure 1, demonstrating that global synchronization ($k_{crit}\to 0$ is achieved for distribution multipliers lower than around $10$, at which point global synchronization causes a uniform drift in the agent ensemble. Critical couplings strengths then increase logarithmically with frequency distribution $\omega_{i}$.
@@ IV Results
The wavelength associated with the Turing-instability analysis (Figure 2a) is indicated with dashed, black line in the first column of Figure 3, and we observe that it roughly coincides with the dominant wavenumber in the structure, providing clear empirical evidence for the analytical derivation of a kinetic Turing instability.
The fast-oscillating modes of Figure 3 start in a chaotic, symmetric state (top row). Symmetry breaks with the triggering of the kinetic Turing instability, and the system is subsequently organized in islands with coherent phase evolution, forming local, stable vortices. These vortices must be understood in terms of the quantized loop currents predicted by Ref. [33], characterized by the sum of phase differences around closed loop in the network being equal to an integer multiple of $2\pi$. Figure 4 demonstrates this, showing the instantaneous velocity of some 2000 agents undergoing clustering. We observe that tracing a circle around the vortices will cycle through all the phases from 0 to $2\pi$, with a topological winding number of -1 (a counter-clockwise vortex) [33]. We also note that (1) the general motion of guiding centers is perpendicular to the motion of the oscillators, and (2) the attraction (guiding centers) is considerably slower than the inhibition (agent motility), providing clear justification for the derived Turing instability.
The characteristic spatial organization that we see in Figures 3 and 4 are robust in the simulations of fast-mode oscillators, suggesting that the formation of quantized loop phase currents may be a universal quality in coupled chiral active matter subject to synchronization rules.
@@
Figure 5 demonstrates that the emergence of structure in the agent ensemble reliably occur for base coupling strengths that exceed the critical value $A_{crit}\approx 0.06$ for the kinetic Turing instability, determined after solving Eqs. (23, 24) for an $\omega_{i}$ distribution multiplier of $100$. This is heralded by the power spectral density of the spatial distribution of active agents (red) and guiding centers (blue) to settle on power law behaviors.
@@ V Discussion
We have simulated ensembles of polar chiral active matter agents that are subject to phase synchronization through a Kuramoto-Sakaguchi-like phase interaction, demonstrating that complex structures emerge from a kinetic Turing instability triggered in the relatively simple interactions between locally coupled oscillators. Through the rigorous derivation of a continuum field theory [28] and empirical evidence (Figure 5), we have demonstrated the crucial role played by the kinetic Turing instability within the chiral active matter ensemble. This configuration builds on the noise-multiplicative self-organization framework of Ref. [3] and is consistent with the theoretical foundation by Ref. [33] in that stable, quantized loop currents of phase information emerge in the ensemble.
Unlike Ref. [17], where the prerequisites for a Turing instability in a fixed oscillator network were coded into a higher-order Kuramoto interaction, we have demonstrated that a kinetic Turing Turing instability can trigger in the spatial distribution of chiral active matter. Our model relies on a localized conventional Kuramoto interaction, where the attractive force mediated by the kernel $G$ creates a positive feedback loop that aligns phase-synchronized agents. Opposing this, the agents’ swim speed $v_{0}$ coupled with their intrinsic frequency dispersion $\omega_{i}$ acts as an effective diffusion mechanism, transporting phase information away from ordered regions and breaking down order. This creates a competition between local activation and kinetic inhibition, driving the bifurcation observed in the linear stability analysis and triggering the instability.
@@
A crucial ingredient in our model is the heterogenity in $\omega_{i}$ [3], the powerlaw distribution in Figure 1. We have confirmed, through rigorous testing, that varying the shape (powerlaw) of this distribution does not substantially change the simulation’s empirical outcome, and we have verified that the stability analysis is not sensitive to which particular powerlaw we select. Nevertheless, it remains for future efforts to map out the specific mechanism governing the non-linear transfer of energy in the system, replacing the semi-analytical approach in Section III. In this regard, a related paper, Ref. [34], derives the hydrodynamic limit of our model, and rigorously demonstrates that widening the distribution in $\omega_{i}$ injects sufficient enstrophy to drive an inertial cascade [3] that eventually culminates in inviscid Euler turbulence.
@@
Our model is, however, constricted to a class of active polar chiral matter, and though the derivation in Section III explicitly bridges kinetic plasma physics and active matter turbulence modeling, future efforts should explicitly address whether self-organized turbulence can enter into the present reductionist explanations for geophysical and astrophysical turbulent structuring [3].
@@ VI Conclusion
We have modeled chiral active matter turbulence by coupling self-propelled agents locally with a Kuramoto-like interaction, endowing the agents with an intrinsic and heterogenic frustration, following Ref. [3]. We have simulated the ensemble, demonstrating that stable spectral scaling laws emerge in the resulting active matter clustering. To explain the results, we derived the spatial pattern formation rules in the active matter, identifying the coupling threshold as a kinetic Turing instability [28]. Following the kinetic theory framework established for active fluids [51, 52, 42], we semi-analytically determined the critical wavenumber selected by the competition between Kuramoto phase-locking and active chiral transport. This provides a rigorous microscopic derivation for the emergent ’chimera-like’ patterns often studied in discrete topologies [17, 24], unifying them under a single continuum instability criterion.